OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 2i OPTIONS: A) f(x) = x^4 - 4x^3 + 10x^2 + 20x + 75 B)f(x) = x^4 - 14x^2 - 40x - 75 C)f(x) = x^4 - 4x^3 - 10x^2 - 20x - 75 D)f(x) = x^4 + 10x^2 - 40x - 75

OpenStudy (anonymous):

see that one zero is -1+2i hence its conjugate that is -1-2i must also be the zero of same polynomial hence required polynomial can be calculated using (x-5)(x+3)(x+1-2i)(x+1+2i)

OpenStudy (anonymous):

i multiplied them but it didn't come out as any of my options… @matricked

OpenStudy (anonymous):

check if you get D

OpenStudy (anonymous):

=(x^2 - 2x -15 )((x+1)^2 - 4i^2) =(x^2 - 2x -15 )(x^2+2x+1+4) as i^2=-1 rest you can multiply...

OpenStudy (anonymous):

typo error for -15

OpenStudy (anonymous):

Oh so it is D! Thanks!

OpenStudy (anonymous):

yw